A researcher has used a one-way analysis of variance model to test whether the average starting salaries differ among recent graduates from the nursing, engineering, business, and education disciplines. She has randomly selected four graduates from each of the four areas. If MSE = 4, and SST= 120, is there a significant difference in the starting salaries among the four disciplines? What are Fstat and Fcrit values?

Select one:

a. No, 6.0, 3.49

b. No, 2.0, 2.49

c. None of the answers

d. Yes, 6.0, 3.49

e. Yes, 2.0, 2.49
Answer #1

Given, MSE = 4 and SST = 120

Let number of groups = k

k = 4 ( as four areas are considered )

Let Total number of items in these groups = n

Here, n = 16 ( as four graduates from each of group are considered )

Fstat = MST / MSE

MST = SST / ( k -1 )

MST = 120 / (4-1) = 120/3 = 40

Therefore, Fstat = 40/4 = 10

Fstat = 10

Fcrit value is observed from F-distribution table. As 'alpha' value is not given let us assume level of significance i.e., alpha at 5%.

F-distribution table at 5% level of significance is shown in the image below:

Appendix 4a 5 per cent Points of the F-distribution Column represents degrees of freedom (v) for numerator of F-test Row repr

From the above table, we can see on the ( n - k )th i.e., 12th ( 16 - 4 ) column and ( k - 1)rd i.e., 3rd ( 4 - 1 ) row we find that Fcrit value which is 3.49.

Fcrit = 3.49

As Fstat (10 ) is higher than Fcrit ( 3.49 ), Yes, there is significant difference between the starting salaries among the four disciplines.

So, the answer to the question is (c).